Question #1

On March 1, 2020, XYZ Company purchased a machine for $150,000. The machine was expected to have a residual value of $24,000 with a useful life of five years or 32,000 units.

Required:

Complete the following table using the double-declining balance method (5 marks)

Year

Depreciation Expense

Book Value

2020

2021

2022

2023

2024
Answer #1

Correct Answer:

Double declining Method

Year

Depreciation expense

Ending Book Value

1

$           50,000

$      100,000

2

$           40,000

$         60,000

3

$           24,000

$         36,000

4

$           12,000

$         24,000

5

$                    -  

$         24,000

Working:

Double declining Method

A

Cost

$          150,000.00

B

Residual Value

$            24,000.00

C=A - B

Depreciable base

$          126,000.00

D

Life [in years]

                                5

E=C/D

Annual SLM depreciation

$            25,200.00

F=E/C

SLM Rate

20%

G=F x 2

DDB Rate

40%

Note: depreciation for the year 2020 will be calculated for 10 months ( 1 March to 31 December)

Depreciation expense for the year 2020 = 150,000 * 40% = 60,000

60,000 / 12 *10

= 5000 * 10

Depreciation Expense for 10 months = $ 50,000

Depreciation Expense for the year 2023 will be calculated by subtracting the residual value from the beginning book value for the year 2023, so that the ending value of the machine do not fall below the residual value.

In the year 2024 , no depreciation will be charged to the machine, as the book value is already equal to the residual value .

Double declining Method

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$            150,000

40%

$           50,000

$      100,000

$       50,000

2

$            100,000

40%

$           40,000

$         60,000

$       90,000

3

$              60,000

40%

$           24,000

$         36,000

$     114,000

4

$              36,000

40%

$           12,000

$         24,000

$     126,000

5

$              24,000

40%

$                    -  

$         24,000

$     126,000

End of Answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks